# JambaTalk: Speech-driven 3D Talking Head Generation based on a Hybrid Transformer-Mamba Model

### v2

@@ Abstract.
In recent years, the talking head generation has become a focal point for researchers. Considerable effort is being made to refine lip-sync motion, capture expressive facial expressions, generate natural head poses, and achieve high-quality video. However, no single model has yet achieved equivalence across all quantitative and qualitative metrics. We introduce Jamba, a hybrid Transformer-Mamba model, to animate a 3D face. Mamba, a pioneering Structured State Space Model (SSM) architecture, was developed to overcome the limitations of conventional Transformer architectures, particularly in handling long sequences. This challenge has constrained traditional models. Jamba combines the advantages of both the Transformer and Mamba approaches, offering a comprehensive solution. Based on the foundational Jamba block, we present JambaTalk to enhance motion variety and lip sync through multimodal integration. Extensive experiments reveal that our method achieves performance comparable or superior to state-of-the-art models. Supplementary video and code are available at: https://JambaTalk.github.io/.
@@ 3. Proposed Method
where represents the feature sequence extracted from the text decoder, and maps the alignment sequence to audio transcription . A projection layer is first used to map the lip feature sequence to the output probabilities, where , , and is the output vocabulary size with a dimension of (Peng et al., 2023a).
@@ 3.3. Transformer Layer
Different rotation speeds (frequencies) are assigned to each 2D subspace, enabling the model to capture both short- and long-range positional information. Each input vector at position is first transformed into two different representations called query and key by multiplying with learned projection matrices and , respectively. These matrices are trained during model learning to map the input features into spaces suitable for attention computation. Next, the Rotary Positional Embedding (RoPE) is applied by multiplying these query and key vectors with a position-dependent rotation matrix for queries and for keys. This rotation matrix encodes positional information by rotating each subspace of the vector based on the token position and predefined frequencies, effectively injecting relative position information into the vectors:
@@
where incorporates relative positional information through the matrix depends only on the position difference . This multiplicative encoding preserves stability and allows the model to capture positional relationships more naturally compared to additive embeddings (Su et al., 2024).
@@
then mapping back to the original dimension:
In these last equations, is a learnable matrix of size that projects the original -dimensional vector into an -dimensional bottleneck, is the compressed representation of the positional embedding (size ), and indicates that is much smaller than , thus greatly reducing parameters count and computational cost. It captures the most important features of the positional embedding in a smaller space to reduce parameters and allow learning. is a learnable matrix that maps the compressed back to the original -dimensional space, and is the final transformed positional embedding (size ). This gives the model a flexible and learnable modification of the original positional embedding, while still keeping the same dimensionality as the original input. The two matrices and form a low-rank approximation of a learnable transformation, which reduces the number of parameters (compared to a full matrix), makes the positional embeddings learnable and more expressive, and keeps the original dimension compatible with downstream layers.
@@
The rotary transformation is then applied to each 2D subspace of the token vector :
This formulation preserves the RoPE's ability to encode relative positions, while allowing task-specific rotation adjustments. The low-rank factorization constrains the additional parameters to , providing flexibility with minimal computational cost. The complete mechanism is summarized in Algorithm 1.
@@ 4.1. Experiments Setting
We employed the publicly accessible 3D dataset, VOCASET (Cudeiro et al., 2019), for training and testing. This dataset includes 480 speech sequences derived from 3D face scans, showcasing the 12 identities found in the CoMA dataset (Ranjan, 2018). The dataset consists of 480 facial motion sequences from 12 subjects. Each sequence is recorded at 60 fps and lasts between 3 and 4 seconds. Each 3D face mesh comprises 5,023 vertices. To provide a fair comparison, we employ the same training, validation, and testing splits as VOCA (Cudeiro et al., 2019), referred to as VOCA-Train, VOCA-Val, and VOCA-Test, respectively. The dataset provides paired audio and 3D facial scans of English utterances. VOCASET includes 255 unique sentences, with some sentences shared across different speakers.
@@ 4.3. Qualitative Evaluation
[…]
Due to GPU limitations (ASUS NVIDIA GeForce RTX 4090), JambaTalk was able to generate predicted animations with input audio of up to 60 seconds. In contrast, Transformer-based competitors such as FaceFormer (Fan et al., 2022) and CodeTalker (Xing et al., 2023) were unable to handle sequences of this length. While SelfTalk (Peng et al., 2023a) could produce an output, its quality was noticeably lower than JambaTalk's, as demonstrated in the published video (https://JambaTalk.github.io/). In particular, SelfTalk (Peng et al., 2023a) exhibits jitter during silent segments, whereas JambaTalk maintains smooth and consistent performance.

### v3

@@ Abstract.
In recent years, the talking head generation has become a focal point for researchers. Considerable effort is being made to refine lip-sync motion, capture expressive facial expressions, generate natural head poses, and achieve high-quality video. However, no single model has yet achieved equivalence across all quantitative and qualitative metrics. We introduce JambaTalk, a hybrid Transformer-Mamba model, to animate a 3D face. Mamba, a pioneering Structured State Space Model (SSM) architecture, was developed to overcome the limitations of conventional Transformer architectures, particularly in handling long sequences. This challenge has constrained traditional models. Jamba combines the advantages of both the Transformer and Mamba approaches, offering a comprehensive solution. Based on the foundational Jamba block, we present JambaTalk to enhance motion variety and lip sync through multimodal integration. Extensive experiments reveal that our method achieves performance comparable or superior to state-of-the-art models. Supplementary video and code are available at: https://JambaTalk.github.io/.
@@ 3. Proposed Method
where represents the sequence of lip features extracted from the text decoder, is the target transcription, is a possible alignment sequence between and . The operator removes repeated labels and blanks, and represents all valid alignment sequences corresponding to .
To compute the CTC loss, a projection layer maps the lip features to output probabilities:
where and are learnable parameters, is the dimension of the lip feature vector , and is the output vocabulary size (Peng et al., 2023a). The predicted probability vector at each time step is then used in the CTC loss to account for all possible alignments between lip features and the target transcription.
@@ 3.3. Transformer Layer
Different rotation speeds (frequencies) are assigned to each subspace, enabling the model to capture both short- and long-range positional information. Each input vector at position is first transformed into two different representations called query and key by multiplying with learned projection matrices and , respectively. These matrices are trained during model learning to map the input features into spaces suitable for attention computation. Next, the Rotary Positional Embedding (RoPE) is applied by multiplying these query and key vectors with a position-dependent rotation matrix for queries and for keys. This rotation matrix encodes positional information by rotating each subspace of the vector based on the token position and predefined frequencies, effectively injecting relative position information into the vectors:
@@
where incorporates relative positional information through the matrix depending only on the position difference . This multiplicative encoding preserves stability and allows the model to capture positional relationships more naturally compared to additive embeddings (Su et al., 2024).
@@
Then, mapping back to the original dimension:
In these last equations, is a learnable matrix whose transpose, , projects the original -dimensional vector into an -dimensional bottleneck, , where . This low-dimensional representation captures the most important features of the positional embedding, reducing the number of parameters and computational cost while allowing learning. Similarly, is a learnable matrix whose transpose, , maps back to the original -dimensional space, producing the final transformed positional embedding . The two matrices and together form a low-rank approximation of a learnable transformation, which makes the positional embeddings more expressive and compatible with downstream layers while keeping the original dimensionality.
@@
The rotary transformation is then applied to each 2 subspace of the token vector :
This formulation preserves the RoPE's ability to encode relative positions, while allowing task-specific rotation adjustments. The low-rank factorization constrains the additional parameters to , providing flexibility at minimal computational cost. The complete mechanism is summarized in Algorithm 1.
@@ 4.1. Experiments Setting
We employed the publicly accessible 3D dataset, VOCASET (Cudeiro et al., 2019), for training and testing. This dataset includes 480 speech sequences derived from 3D face scans, showcasing the 12 identities found in the CoMA dataset (Ranjan, 2018). Each 3D face mesh comprises 5,023 vertices. To provide a fair comparison, we employ the same training, validation, and testing splits as VOCA (Cudeiro et al., 2019), referred to as VOCA-Train, VOCA-Val, and VOCA-Test, respectively. The dataset provides paired audio and 3D facial scans of English utterances. VOCASET includes 255 unique sentences, with some sentences shared across different speakers.
@@ 4.3. Qualitative Evaluation
On a single NVIDIA GeForce RTX 4090 (24 GB), JambaTalk can process audio-driven sequences up to 60 seconds in a single pass. Transformer-based baselines such as FaceFormer (Fan et al., 2022) and CodeTalker (Xing et al., 2023) cannot process sequences of this length due to the quadratic memory cost of attention mechanisms. SelfTalk (Peng et al., 2023a) can handle long sequences similar to JambaTalk but exhibits noticeable quality degradation, including jitter during silent segments, whereas JambaTalk maintains stable and smooth performance (see supplementary video at https://JambaTalk.github.io/).
Diffusion-based methods such as DiffPoseTalk (Sun et al., 2024) can handle sequences longer than those demonstrated in our experiments (up to 2 minutes), but rely on multi-step denoising and sliding-window schemes that involve different computational trade-offs. Our work specifically targets improving the long-sequence capacity of Transformer-style models through the linear-complexity Mamba component.
@@ 5. Ablation Study
| Metric | Range (Typical) | Interpretation |
| Inference Time (Latency) | 30.6 ms – 42.7 ms | Time required to process a 1-second audio window; low latency suitable for real-time applications. |
| Output FPS (Throughput) | 600 – 800 FPS | Rate at which the model generates mesh frames. |
| Speedup (vs. real-time audio) | 24 – 32 | Frames generated roughly 30× faster than real-time audio playback. |
6. Runtime Efficiency Analysis
We have implemented a real-time, speech-driven 3D talking face generation pipeline that efficiently handles both microphone and audio file inputs. Audio signals are streamed into a queue in small chunks, enabling continuous processing without blocking the user interface. Each chunk is converted into features using a Wav2Vec2-based processor and cast to half precision (float16) to optimize GPU memory usage. The model processes these features in overlapping windows, allowing low-latency inference while maintaining temporal continuity. This approach supports high-throughput frame generation, with each audio window producing the corresponding 3D mesh vertices for the talking head.
To render the generated sequences, a dedicated off-screen 3D renderer updates the mesh in real time, providing smooth and interactive visualization. Mesh predictions are placed into a secondary queue consumed by the rendering thread, ensuring synchronization between inference and display. This design enables efficient pipelining, allowing JambaTalk to achieve real-time performance on standard GPUs (e.g., NVIDIA RTX 4090) with low latency and significant speedup over real-time audio playback. The system can generate infinite-length sequences using a sliding-window approach. Runtime performance metrics are summarized in Table 7. (see supplementary video at https://JambaTalk.github.io/).
@@ Acknowledgements.
- Sun et al. (2024) Zhiyao Sun, Tian Lv, Sheng Ye, Matthieu Lin, Jenny Sheng, Yu-Hui Wen,and Yong-jin Liu. 2024. Diffposetalk: Speech-driven stylistic 3d facial animation and head pose generation via diffusion models. CM Transactions on Graphics (TOG) (2024).